# StableVC: Style Controllable Zero-Shot Voice Conversion with Conditional Flow Matching

### v1

@@ Abstract
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to an arbitrary unseen speaker while preserving the original linguistic content. Despite recent advancements in zero-shot VC using language model-based or diffusion-based approaches, several challenges remain: 1) current approaches primarily focus on adapting timbre from unseen speakers and are unable to transfer style and timbre to different unseen speakers independently; 2) these approaches often suffer from slower inference speeds due to the autoregressive modeling methods or the need for numerous sampling steps; 3) the quality and similarity of the converted samples are still not fully satisfactory. To address these challenges, we propose a Style controllable zero-shot VC approach named StableVC, which aims to transfer timbre and style from source speech to different unseen target speakers. Specifically, we decompose speech into linguistic content, timbre, and style, and then employ a conditional flow matching module to reconstruct the high-quality mel-spectrogram based on these decomposed features. To effectively capture timbre and style in a zero-shot manner, we introduce a novel dual attention mechanism with an adaptive gate, rather than using conventional feature concatenation. With this non-autoregressive design, StableVC can efficiently capture the intricate timbre and style from different unseen speakers and generate high-quality speech significantly faster than real-time. Experiments demonstrate that our proposed StableVC outperforms state-of-the-art baseline systems in zero-shot VC and achieves flexible control over timbre and style from different unseen speakers. Moreover, StableVC offers approximately 25$\times$ and 1.65$\times$ faster sampling compared to autoregressive and diffusion-based baselines. Audio samples can be found in demo pages  [^1] https://yaoxunji.github.io/stablevc/.
@@ Introduction
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to an arbitrary unseen speaker while keeping the linguistic content unchanged. This technology has seen significant advancements, finding crucial applications in both professional audiobook production and entertainment short videos (Ning et al. 2023b; Yao et al. 2023). Currently, numerous impressive zero-shot VC techniques demonstrate remarkable efficacy in converting realistic and natural sound samples (Hussain et al. 2023; Wang et al. 2023b; Li et al. 2024; Popov et al. 2021; Choi, Lee, and Lee 2024). However, these approaches primarily focus on adapting timbre to unseen speakers while often overlooking style attributes. Style refers to how a person expresses themselves, including pitch variation, speaking rate, intonation, and emotion. It can vary significantly across different contexts to enhance speech expressiveness. Additionally, most current approaches come at a cost, such as more complicated training setups, additional quantization steps, and computationally expensive autoregressive formulations.
@@
To achieve style-controllable VC and tackle inference speed problems, we propose StableVC, a style-controllable zero-shot VC model that is fast and efficient. StableVC first disentangles speech into content, timbre, and style, and then employs a flow matching generative module to reconstruct high-quality mel-spectrograms. The flow matching module consists of multiple diffusion transformer (DiT) blocks and is trained to predict a vector field, which efficiently models the probabilistic distribution of the target mel-spectrograms. For attribute disentanglement, we use a pre-trained self-supervised model with K-means clustering to extract linguistic content, while a factorized codec is used to extract style representation. Instead of using conventional concatenation for timbre and style modeling, we propose a novel dual attention mechanism with an adaptive gate control to capture timbre and style information effectively. An adaptive style gate and strong timbre prior information are introduced to ensure the stability of timbre and style modeling. Therefore, we can convert timbre to the target unseen speaker while flexibly controlling style using another speaker’s reference. The main contributions of this study can be summarized as follows:
@@
A popular paradigm for zero-shot VC involves tokenizing speech waveforms into discrete semantic and acoustic tokens using a self-supervised learning (SSL) model and a neural audio codec, respectively. This approach allows the generation task to leverage powerful large language models, exhibiting impressive zero-shot capabilities. For example, LM-VC (Wang et al. 2023b) employs a two-stage language modeling approach: first generating coarse acoustic tokens to recover the source linguistic content and target speaker’s timbre, and then reconstructing fine acoustic details as converted speech. Similarly, several zero-shot VC frameworks (Yang et al. 2024; Li et al. 2024; Hussain et al. 2023; Wang et al. 2022) use semantic representations extracted from SSL models to disentangle the linguistic content. GR0 (Wang et al. 2024) uses a generative SSL framework to jointly learn a global speaker embedding and a zero-shot voice converter, achieving zero-shot VC. Another approach (Luo and Dixon 2024) decouples speech into local and global representations and employs a dropout technique and multiplicative Gaussian noise to apply an information bottleneck for timbre disentanglement. Although these approaches achieve zero-shot voice conversion, they mainly focus on capturing the timbre and overlook the style information.
@@
Our study aims to provide flexible style control capabilities in zero-shot VC with high quality and efficiency. We begin with an overview of our approach, followed by a detailed explanation of the proposed DualAGC mechanism. Next, we describe the flow matching module and conclude with a discussion of the overall training objectives.
@@
The overall framework of StableVC is illustrated in Figure 2. First, we extract linguistic content, style representation, and mel-spectrogram from the waveform. To disentangle linguistic content, we use a pre-trained WavLM model [^2] https://github.com/microsoft/unilm/tree/master/wavlm and apply K-means clustering to extract discrete tokens (Chen et al. 2022), setting the number of K-means clusters to 1024. Additionally, we deduplicate adjacent tokens and replace them with the corresponding K-means embeddings. This process is the same as described in Yao et al. 2024 and has demonstrated its effectiveness in linguistic content extraction. The purpose of deduplication is to re-predict the duration for each token, allowing the duration of the same linguistic content to adapt to different timbre and styles. For style representation, we use a factorized codec [^3] https://github.com/open-mmlab/Amphion/tree/main/models/codec/ns3˙codec as style extractor. This model can extract disentangled subspace representations of speech style. For timbre representation, we extract the mel-spectrogram from multiple reference speeches by the same speaker as the source speech. The details of timbre and style modeling will be discussed in the following section.
@@
Subjective Metrics: We employ the naturalness mean opinion score (nMOS) to evaluate the naturalness of the generated samples. Additionally, we use two similarity mean opinion scores (sMOS-s and sMOS-p) to evaluate style similarity and timbre similarity, respectively. We invite 15 professional participants to listen to the generated samples and provide their subjective perception scores.

### v2

@@ Abstract
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to an arbitrary unseen speaker while preserving the original linguistic content. Despite recent advancements in zero-shot VC using language model-based or diffusion-based approaches, several challenges remain: 1) current approaches primarily focus on adapting timbre from unseen speakers and are unable to transfer style and timbre to different unseen speakers independently; 2) these approaches often suffer from slower inference speeds due to the autoregressive modeling methods or the need for numerous sampling steps; 3) the quality and similarity of the converted samples are still not fully satisfactory. To address these challenges, we propose a Style controllable zero-shot VC approach named StableVC, which aims to transfer timbre and style from source speech to different unseen target speakers. Specifically, we decompose speech into linguistic content, timbre, and style, and then employ a conditional flow matching module to reconstruct the high-quality mel-spectrogram based on these decomposed features. To effectively capture timbre and style in a zero-shot manner, we introduce a novel dual attention mechanism with an adaptive gate, rather than using conventional feature concatenation. With this non-autoregressive design, StableVC can efficiently capture the intricate timbre and style from different unseen speakers and generate high-quality speech significantly faster than real-time. Experiments demonstrate that our proposed StableVC outperforms state-of-the-art baseline systems in zero-shot VC and achieves flexible control over timbre and style from different unseen speakers. Moreover, StableVC offers approximately 25$\times$ and 1.65$\times$ faster sampling compared to autoregressive and diffusion-based baselines.
@@ Introduction
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to an arbitrary unseen speaker while keeping the linguistic content unchanged. Currently, numerous impressive zero-shot VC techniques demonstrate remarkable efficacy in converting realistic and natural sound samples, finding crucial applications in both professional audiobook production and entertainment short videos (Popov et al. 2021; Ning et al. 2023b; Yao et al. 2023; Li et al. 2024; Choi, Lee, and Lee 2024). However, these approaches primarily focus on adapting timbre to unseen speakers while often overlooking style attributes. Style refers to how a person expresses themselves, including pitch variation, speaking rate, intonation, and emotion. It can vary significantly across different contexts to enhance speech expressiveness. Additionally, most current approaches come at a cost, such as more complicated training setups and computationally expensive autoregressive formulations.
@@
To achieve style-controllable VC and tackle inference speed problems, we propose StableVC, a style-controllable zero-shot VC model that is fast and efficient. StableVC first disentangles speech into content, timbre, and style, and then employs a flow matching generative module to reconstruct high-quality mel-spectrograms. The flow matching module consists of multiple diffusion transformer (DiT) blocks and is trained to predict a vector field, which efficiently models the probabilistic distribution of the target mel-spectrograms. For attribute disentanglement, we use a pre-trained self-supervised model with K-means clustering to extract linguistic content, while a factorized codec is used to extract style representation. Instead of using conventional concatenation for timbre and style modeling, we propose a novel dual attention mechanism with an adaptive gate control to capture timbre and style information effectively. An adaptive style gate and strong timbre prior information are introduced to ensure the stability of timbre and style modeling. Therefore, we can convert timbre to the target unseen speaker while flexibly controlling style using another speaker’s reference. Audio samples can be found in demo pages  [^1] https://yaoxunji.github.io/stablevc/. The main contributions of this study can be summarized:
@@
A popular paradigm for zero-shot VC involves tokenizing speech waveforms into discrete semantic and acoustic tokens using a self-supervised learning (SSL) model and a neural audio codec, respectively. This approach allows the generation task to leverage powerful large language models, exhibiting impressive zero-shot capabilities. For example, LM-VC (Wang et al. 2023b) employs a two-stage language modeling approach: first generating coarse acoustic tokens to recover the source linguistic content and target speaker’s timbre, and then reconstructing fine acoustic details as converted speech. Similarly, several zero-shot VC frameworks (Yang et al. 2024; Li et al. 2024; Hussain et al. 2023; Wang et al. 2022) use semantic representations extracted from SSL models to disentangle the linguistic content. GR0 (Wang et al. 2024) uses a generative SSL framework to jointly learn a global speaker embedding and a zero-shot voice converter, achieving zero-shot VC. Another approach (Luo and Dixon 2024) decouples speech into local and global representations and employs dropout with multiplicative Gaussian noise to apply an information bottleneck for timbre disentanglement. Although these approaches achieve zero-shot VC, they mainly focus on capturing the timbre and overlook the style information.
@@
We aim to provide flexible style control capabilities in zero-shot VC with high quality and efficiency. The overall framework of StableVC is illustrated in Figure 2. First, we extract linguistic content, style representation, and mel-spectrogram from the waveform. To disentangle linguistic content, we use a pre-trained WavLM model [^2] https://github.com/microsoft/unilm/tree/master/wavlm and apply K-means clustering to extract discrete tokens (Chen et al. 2022), setting the number of K-means clusters to 1024. Additionally, we deduplicate adjacent tokens and replace them with the corresponding K-means embeddings. This process is the same as described in Yao et al. 2024 and has demonstrated its effectiveness in linguistic content extraction. The purpose of deduplication is to re-predict the duration for each token, allowing the duration of the same linguistic content to adapt to different timbre and styles. For style representation, we use a factorized codec [^3] https://github.com/open-mmlab/Amphion/tree/main/models/codec/ns3˙codec as style extractor. This model can extract disentangled subspace representations of speech style. For timbre representation, we extract the mel-spectrogram from multiple reference speeches by the same speaker as the source speech. The details of timbre and style modeling will be discussed in the following section.
@@
Subjective Metrics: We employ the naturalness mean opinion score (nMOS) to evaluate the naturalness of the generated samples. Additionally, we use two similarity mean opinion scores (sMOS-s and sMOS-p) to evaluate style similarity and timbre similarity, respectively.